### v1

# Advancing Embodied Intelligence in Robotic-Assisted Endovascular Procedures: A Systematic Review of AI Solutions

Abstract
Endovascular procedures have revolutionized the treatment of vascular diseases thanks to minimally invasive solutions that significantly reduce patient recovery time and enhance clinical outcomes. However, the precision and dexterity required during these procedures poses considerable challenges for interventionists. Robotic systems have emerged offering transformative solutions, addressing issues such as operator fatigue, radiation exposure, and the inherent limitations of human precision. The integration of Embodied Intelligence (EI) into these systems signifies a paradigm shift, enabling robots to navigate complex vascular networks and adapt to dynamic physiological conditions. Data-driven approaches, advanced computer vision, medical image analysis, and machine learning techniques, are at the forefront of this evolution. These methods augment procedural intelligence by facilitating real-time vessel segmentation, device tracking, and anatomical landmark detection. Reinforcement learning and imitation learning further refine navigation strategies and replicate experts’ techniques. This review systematically examines the integration of EI principles into robotic technologies, in relation to endovascular procedures. We discuss recent advancements in intelligent perception and data-driven control, and their practical applications in robot-assisted endovascular procedures. By critically evaluating current limitations and emerging opportunities, this review establishes a framework for future developments, emphasizing the potential for greater autonomy and improved clinical outcomes. Emerging trends and specific areas of research, such as federated learning for medical data sharing, explainable AI for clinical decision support, and advanced human-robot collaboration paradigms, are also explored, offering insights into the future direction of this rapidly evolving field.
Index Terms:
robotic-assisted endovascular procedures, medical image analysis, reinforcement learning
I Introduction
Endovascular procedures have been subject to a paradigm shift in relation to the management of vascular disease with the advent of minimally invasive therapeutic strategies that significantly shorten patient recovery times and enhance clinical outcomes [1]. Central to these procedures is the intricate manipulation of catheters and guidewires under real-time imaging guidance, a task that necessitates exceptional technical skill and dexterity, as illustrated in Figure 1(a), which outlines the workflow of an endovascular procedure [2]. Despite unparalleled progress, conventional manual approaches are constrained by factors such as operator fatigue, radiation exposure, and the inherent shortcomings of human precision [3, 4]. The emergence of robotic systems has begun to address these challenges, providing enhanced stability, accuracy, and reproducibility, thereby augmenting the efficacy of these procedures [5]. Nevertheless, current robotic systems still use leader-follower technologies, as depicted in Figure 1(b), in which the operator manipulates the control console from within the operational cockpit. In this configuration, the robotic system translates the operator’s inputs into precise movements of the catheter and guidewire, facilitating remote control and minimizing the need for direct physical interaction within the patient.
Figure 1: Overview of an endovascular procedure. Instruments, such as catheters and guidewires, are inserted into and navigated through the patient’s vascular system to reach the target site for treatment. (b) Schematic representation of a robotic-assisted endovascular procedure. The primary components of the robotic system include the operator situated at the operational cockpit, controlling the leader console. This console transmits commands to the follower robotic drive, which manipulates the instruments inside the patient. Real-time imaging is provided by the C-arm X-ray machine, guiding the procedure with the patient on the operating table. (c) Conceptual framework for embodied intelligence (EI) in Robotic-Assisted Endovascular Procedures. AI agents facilitate task-level autonomy through reinforcement learning (RL), task modeling, and real-time data analysis. By interacting with both simulated and real-world environments, the system adapts and optimizes procedural strategies. Core elements of perception, interaction, and decision-making are supported by vascular anatomy models and patient-specific digital twin simulations. (d) Composition of EI in robotic-assisted endovascular procedures. The system employs advanced perception techniques for vascular segmentation and device tracking to ensure precise navigation of interventional tools. AI-driven control strategies continuously refine task execution through learning-based approaches. Key challenges encompass technical limitations, such as real-time data integration, alongside ethical and clinical considerations regarding patient safety and regulatory compliance. The schematic was created using BioRender (https://biorender.com).
Alongside this evolving landscape, the concept of Embodied Intelligence (EI) has emerged as a pivotal theoretical and technological driver [6, 7], particularly within the medical domain [8, 9]. EI diverges from traditional robotic design philosophies by emphasizing the interplay between a robot’s physical embodiment and its intelligent behaviors [10]. This is particularly relevant for endovascular robotics, in which systems must navigate complex, deformable vascular networks and be able to adapt, in real-time, to a continuously shifting physiological environment [11]. By incorporating EI principles, future robotic platforms can transcend pre-defined, scripted responses, and begin moving towards adaptive, contextually aware capabilities that more closely mimic the nuanced skills of experienced cardiologists [12, 13]. As depicted in Figure 1(c), EI can also serve as a co-pilot, assisting cardiologists by providing real-time decision support and optimizing procedural strategies.
The integration of data-driven techniques provides the computational foundation for realizing Embodied Intelligence (EI) in robotic-assisted endovascular procedures. Advanced computational methodologies, from computer vision and machine learning-enhanced medical image analysis to reinforcement and imitation learning, hold the potential to distill, harness and replicate the expertise of experienced clinicians. These methodologies facilitate precise anatomical interpretation, improved device tracking, and adaptive procedural planning, effectively merging the cognitive and motor elements of interventional practice into a unified, intelligent robotic architecture.
Over time, endovascular robotics has evolved from rudimentary platforms offering remote manipulation and limited motion control to sophisticated systems equipped with multi-modal sensing and intelligent decision support. With the advent of Artificial Intelligence (AI) and Machine Learning (ML), next-generation systems now strive for greater autonomy, incorporating automated path planning and real-time obstacle avoidance into their repertoires. Empirical evidence from preclinical studies underscores the promise of these developments, heralding a future in which endovascular procedures are safer, more efficient, and decreasingly operator-dependent [14].
Recent breakthroughs in deep learning and computer vision have further accelerated progress toward integrated perception and control. State-of-the-art algorithms enable real-time vascular segmentation [15] and device localization [16], facilitating on-the-fly adjustments to operative strategies. Meanwhile, the application of reinforcement learning is helping to refine navigation and decision-making policies, allowing robotic agents to adapt autonomously to patient-specific vascular geometries and pathological conditions. Hybrid solutions combining data-driven autonomy with traditional control frameworks now benefit from robust engineering principles and the flexibility afforded by advanced computational intelligence. As depicted in Figure 2, a chord diagram illustrates the intricate interplay between perception and control modules in robotic-assisted endovascular procedures. The thickness of each chord represents the frequency of co-occurrence of these modules in the literature, quantified through an adjacency matrix derived from relevant studies. These frequency values are normalized to ensure visual clarity, with thicker chords indicating higher co-occurrence, highlighting the integrated roles of these modules in enhancing procedural outcomes. The perception modules, encompassing vessel segmentation, instrument detection and tracking, and multi-modal registration, process real-time imaging data to provide spatial awareness and anatomical context. These feed into the control modules, which include autonomous navigation, enabling precise manipulation of interventional tools. Clinical integration serves as a pivotal bridge, facilitating bidirectional communication between perception and control to optimize procedural outcomes. The diagram’s directional arrows and varying line thicknesses highlight the strength and flow of these connections, underscoring their collaborative role in advancing endovascular interventions.
Despite these advancements, significant challenges persist. The inherent variability in patient-specific anatomy and pathology necessitates robust generalization capabilities, while stringent safety and regulatory requirements also have to be met. Limitations in the availability of high-quality, annotated medical datasets and the computational demands of real-time processing further constrain algorithmic performance. Moreover, the integration of novel robotic technologies into established clinical workflows requires careful consideration of regulatory hurdles as well as clinician acceptance, which is ultimately dependent upon demonstrable improvements in patient outcomes.
This comprehensive review examines the state of data-driven endovascular robotics through the lens of EI. Advances in intelligent perception and data-driven control methodologies are discussed, along with an evaluation of current capabilities, key challenges, and translational potential. By critically appraising recent achievements and identifying areas of opportunity, this review provides a conceptual framework that may help guide future innovations. Emerging trends and specific areas of research, such as federated learning for secure and distributed data sharing, explainable AI for transparent clinical decision support, and refined paradigms for human-robot collaboration are discussed. By offering this analysis, we hope to illuminate the potential of EI-enabled endovascular robotics to reshape the horizons of minimally invasive vascular care.
The structure of this review is presented in Figure 1(d). Section II explores intelligent perception techniques in endovascular procedures, emphasizing recent progress in vascular segmentation and intervention device tracking. These advancements are essential for improving the precision of such interventions. Section III explores the integration of machine learning in robotic-assisted endovascular procedures, providing an in-depth analysis of its role in advancing intelligence and refining control for these complex operations. Section IV assesses the major challenges facing the field, encompassing the limited availability of high-quality coronary imaging data, the lack of a standardized evaluation framework, constraints on the scalability of perception and control methodologies, ethical and legal implications, and the complexities involved in incorporating these technologies into clinical workflows.
Figure 2: Chord diagram depicting the interplay between perception and control modules in robotic-assisted endovascular procedures. Perception modules (blue) include vessel segmentation, instrument detection and tracking, and multi-modal registration, linked to control modules (purple) via clinical integration. Control encompasses autonomous navigation. Arrows denote directional relationships, with line thickness reflecting connection strength, illustrating their integrated roles in enhancing procedural outcomes.
II Intelligent Perception Techniques in Endovascular Procedures
Endovascular procedures critically rely on precise and robust scene understanding to ensure effective navigation and task execution [17, 18]. Scene perception in this domain encompasses a range of essential tasks, such as vessel segmentation, instrument detection and tracking, and the recognition of anatomical structures across multiple imaging modalities. Instrument tracking is another key imperative to ensure the accurate placement of devices and mitigate potential complications, particularly in complex vascular regions [19]. Digital Subtraction Angiography (DSA) has become the default imaging technique, on account of its capacity to provide high-resolution, real-time visualization of blood vessels [20]. However, achieving accurate scene understanding remains a significant challenge, primarily due to the high variability in vascular anatomies, noise in imaging data, and the necessity for real-time processing to support dynamic procedural demands.
Figure 3: Examples of perception tasks in endovascular procedures include the following: (a) Intraoperative vessel segmentation from DSA images to obtain detailed vascular anatomical structures [21]. (b) Device detection and localization utilizing intraoperative DSA images, providing essential visual feedback for cardiologists and robotic systems [22]. (c) multi-modal image registration between preoperative CTA and intraoperative DSA to facilitate interventional navigation [23]. The images presented here are reproduced from [21], [22], and [23], respectively.
Recent developments, particularly in deep learning methodologies, have substantially enhanced the capabilities of these systems, ultimately leading to improved clinical efficacy and patient safety [24, 25]. To further refine navigational precision, 2D intraoperative DSA is frequently fused with 3D preoperative Computed Tomography Angiography (CTA) through multi-modal image registration techniques [23]. This registration process is pivotal to the accurate alignment of the two imaging modalities required to provide comprehensive spatial information. By integrating high-resolution, real-time DSA with detailed 3D anatomical information from CTA, cardiologists can achieve superior navigation and enhanced visualization of complex vascular structures, ultimately improving procedural accuracy and clinical outcomes [26].
II-A Vascular Segmentation Approaches
Vascular segmentation algorithms are carefully designed to delineate the vascular network with high precision, enabling accurate localization of target lesions and supporting optimal device deployment. Furthermore, anatomical structure recognition plays a critical role in providing a comprehensive understanding of the procedural environment. By identifying key anatomical features, such as bifurcations, stenosis, and tortuosity, and capturing the intricate connectivity of the vascular topology, these algorithms can enhance navigational capabilities for both cardiologists and robotic systems. This not only augments the accuracy of interventions but also contributes to safer and more efficient procedural work by ensuring improved localization of critical anatomical structures.
Recent advancements in vascular segmentation have demonstrated notable progress through various deep-learning approaches. These advancements can be broadly categorized into three groups: (1) supervised learning, whereby large annotated datasets are employed to train models for precise vessel delineation; (2) weakly-supervised learning, which aims to diminish the dependency on labeled data; and (3) clinical applications, specifically tailored to address the challenges encountered in multi-center and real-world clinical studies.
II-A1 Supervised Learning Methods
Traditional supervised approaches have been shown to be effective in vascular segmentation, particularly through the utilization of the UNet architecture and its various adaptations, as illustrated in Figure 4. These architectures have emerged as the dominant paradigm in coronary artery segmentation due to their ability to capture fine-grained features of vascular structures. Wang et al. proposed a spatial-temporal segmentation framework that effectively addresses motion in angiographic videos [27]. Their method incorporates a motion compensation mechanism alongside multi-frame fusion to preserve temporal consistency and mitigate motion artifacts, achieving an Intersection over Union (IoU) of 0.714. Shi et al. introduced UENet, leveraging adversarial learning for enhanced feature extraction [28]. This GAN-based architecture employs multi-level feature enhancement and a boundary-aware loss function, resulting in a pixel accuracy of 0.8353. Notable improvements were made by Han et al.’s recursive centerline-aware network [29], which integrated directional awareness and ensemble learning to achieve better vessel delineation, and an F1 score of 0.831.
Other notable contributions in supervised learning include Hao et al.’s attention-guided sequential learning framework [30], which achieved an F1 score of 0.843 by incorporating a channel-wise attention mechanism and ensuring frame-to-frame consistency. Li et al. developed CAU-Net with pyramid feature fusion [31], utilizing cascaded attention units and multi-scale pyramid features to attain a Dice score of 0.904. Liu et al. proposed a full-resolution network with dual thresholding [32], maintaining high-resolution features throughout the network and employing adaptive thresholding, while achieving F1 scores of 0.802 and 0.760 on the DCA1 and CHUAC datasets, respectively. Zhang et al. introduced a recursive centerline-aware network that incorporates direction-aware feature learning and ensemble prediction mechanisms [33]. Their approach achieved an F1 score of 0.831, demonstrating the effectiveness of recursive refinement in capturing intricate vessel patterns. Building on this progress, Zhang et al. proposed a novel contrastive diffusion-based learning framework that integrates a UNet architecture with a sophisticated channel attention module and centerline auxiliary supervision [34]. This method achieved a Dice score of 0.8528, highlighting how the combination of attention mechanisms and auxiliary supervision can further enhance segmentation accuracy.
Figure 4: Three representative approaches to weakly-supervised segmentation for vascular structures are presented. (a) A Generative Adversarial Network (GAN)-based architecture that incorporates contrastive learning to leverage unlabeled data for vessel segmentation [20]. (b) A diffusion-based method that progressively refines vessel structures from noisy initializations to achieve precise segmentation results [35]. (c) An endpoint-based framework that features the Coarse-to-Fine Keypoint Detection Network (CFKD-Net) with Multi-branch Feature Aggregation (MFA) [36]. The images are reproduced from [20], [35], and [36], respectively.
Figure 5: State-of-the-art methods for detecting and localizing interventional devices in X-ray fluoroscopy are discussed. (a) A two-stage framework that incorporates temporal information and Hessian-based features for guidewire detection and segmentation has been proposed [37]. (b) A Light UNet architecture utilizing synthetic data and transfer learning has been employed for catheter segmentation [38]. (c) An efficient network architecture utilizing recurrent convolutional layers has been developed for real-time catheter segmentation and tracking [39]. (d) A fast recurrent attention network designed for simultaneous mono- and dual-guidewire segmentation and tracking has been introduced [22]. (e) A deep learning approach based on coordinate regression has been proposed for automated catheter detection in structural heart interventions [40]. The images are reproduced from [37], [38], [39], [22], and [40].
II-A2 Weakly-supervised Learning
To address the challenge posed by limited labeled data, researchers have explored weakly-supervised strategies, particularly through the generation of pseudo-labels. These approaches aim to leverage vast amounts of unlabeled data while minimizing the reliance on costly manual annotations. Zhang et al. developed a weakly supervised framework that utilizes noisy labels and self-paced learning, achieving a Dice score of 0.821 [41]. Their method incorporates curriculum learning and progressive refinement to effectively manage noisy labels. Ma et al. introduced a self-supervised approach that employs adversarial learning without the need for manual labels, reporting a Dice score of 0.557 on the XCAD dataset [20]. Their method leverages label-free training and consistency regularization to learn representative features from unlabeled data. More recently, Wu et al. proposed DeNVeR, an unsupervised deformable neural representation method of video segmentation, demonstrating the potential of label-free learning [42]. This approach achieved a Dice score of 0.733 across multiple datasets (XCAD, CADICA, and XACV) by leveraging deformable modeling and temporal consistency. Kim et al. introduced a contrastive diffusion-based learning approach, achieving Dice scores of 0.661, 0.700, and 0.621 on the XCAD, 134XCA, and 30XCA datasets respectively [35]. Their method utilizes label-free diffusion models and contrastive learning to extract meaningful features from unlabeled data.
II-A3 Advanced Clinical Applications
Unlike research that focuses solely on algorithm optimization, studies aimed at clinical applications emphasize solutions that prioritize reliability, adaptability, and consistency. Factors such as patient variability, noisy imaging data, and the requirement for real-time performance pose unique challenges in this domain. Song et al. developed a width-aware UNet specifically tailored for coronary vessel analysis, achieving a Dice score of 0.773 in clinical validation [43]. Their method incorporates vessel width prediction and an enhanced UNet architecture designed for clinical assessment. Nobre Menezes et al. conducted a comprehensive multicentric validation study, providing valuable insights into the real-world performance of these systems [44]. Their large-scale validation across multiple centers reported an impressive Dice score of 0.948, demonstrating the potential for clinical adoption. Recent innovations include Kim et al.’s contrastive diffusion learning approach [35] and Guo et al.’s edge-aware vessel segmentation method [45], both of which address specific clinical challenges. Kim et al.’s approach leverages contrastive learning and diffusion models to enhance segmentation performance in low-contrast scenarios, whereas Guo et al.’s method focuses on preserving edge information and refining vessel boundaries, achieving a Dice score of 0.734 on the XCAD dataset. Wang et al. introduced CE-NC-VesselSegNet for clinical use, demonstrating the effectiveness of cross-modality and transfer learning strategies [46]. Their method achieved Dice scores of 0.856 and 0.738 on two different clinical cohorts, showcasing its adaptability to various clinical settings. Yang et al. proposed a geometry-constrained cascaded network, integrating geometric priors and multi-stage refinement to achieve Dice scores of 0.778 and 0.895 on the ASOCA and CCA-200 datasets respectively [47].
II-A4 Available Public Datasets
The availability of large-scale, publicly accessible datasets for coronary vascular segmentation remains limited, posing challenges for reproducibility and standardized benchmarking. Due to clinical regulations and privacy concerns, most current research relies on private datasets, further constraining the accessibility of standardized resources for broader evaluation and comparison. However, several valuable public datasets have been developed, as summarized in Table I. The XCAD dataset [20], intended for weakly-supervised learning, comprises 1621 training frames and 126 test angiograms. DCA1 and CHUAC, utilized in multi-dataset validation studies [32], provide 134 expert-annotated X-ray coronary angiograms and 30 diverse angiography images, respectively. The ASOCA CTA dataset [48] contributes 60 CCTA cases with expert annotations, suitable for the development of 3D segmentation algorithms. CCA-200 [47] presents 200 CCTA images annotated by professional radiologists, while CADICA [42] offers over 6,000 invasive coronary angiography images, supporting unsupervised learning approaches.
TABLE I: Publicly Available Datasets for Coronary Vessel Segmentation
Dataset Description
XCAD [20] 1621 training frames (masks and angiograms), 126 test angiograms with expert annotations. Designed for weakly-supervised learning.
DCA1 [49] 134 X-ray coronary angiograms with expert-outlined ground truth.
CHUAC [50] 30 diverse coronary angiography images.
ASOCA CTA [48] 40 training and 20 testing cases of coronary CT angiography (CCTA) images with expert annotations. 30 patients with coronary artery disease.
CCA-200 [47] 200 CCTA images of coronary artery disease with professional radiologist annotations of coronary artery lumen.
CADICA [42] Over 6,000 invasive coronary angiography images. Suitable for unsupervised learning approaches.
Note: This table summarizes publicly available datasets for vascular segmentation research, including their size and content.
II-B Intervention Device Perception
In endovascular procedures, catheters and guidewires can be effectively modeled as continuum robots [51]. Traditional control systems for these flexible robots have often relied on embedded sensors for state estimation, necessitating the integration of sensors within the robot’s body structure. However, the emergence of image data-driven and image-based algorithms presents a promising alternative [52, 53, 54, 55]. These methods enable the estimation of the configuration, shape, and position of flexible instruments without the need for failure-prone embedded sensors [56, 57]. This paradigm shift not only simplifies device design but also enhances overall system robustness by reducing its dependence on complex and often fragile embedded sensors, such as Fiber Bragg Grating (FBG) sensors [58].
Image-based state perception methods leverage readily available imaging data to infer the 3D shape of continuum instruments and provide real-time positional feedback for the end-effector. These techniques often utilize computer vision and machine learning algorithms to extract relevant features from images and reconstruct the instrument’s geometry. Multi-modal data fusion, which integrates information from multiple imaging modalities, can then further enhance the accuracy and completeness of state estimation.
To provide a structured overview of recent advancements, a table summarizing key developments in intelligent perception techniques for endovascular procedures is presented below. The table is organized chronologically from 2020 to 2024 and categorized as follows: Vessel Segmentation, encompassing various algorithms and architectures for the automated delineation of vessel structures from different imaging modalities; Instrument Detection and Tracking, focusing on techniques for real-time localization and tracking of guidewires and catheters within the vascular system; Continuum Robot State Estimation, detailing methods for inferring the 3D configuration, shape, and position of flexible instruments from image data; and multi-modal Registration, describing techniques for aligning and integrating information from multiple imaging modalities, such as DSA and CTA, to provide a comprehensive view of the procedural environment. This structured approach aims to facilitate a deeper understanding of the current state-of-the-art and to highlight future research directions in the field of intelligent perception for endovascular procedures.
TABLE II: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception
No. Authors Category Method and Key Features Performance Metrics Dataset
1 Wang et al. [27] Supervised Learning for coronary vessel segmentation Spatio-temporal CNN with motion compensation mechanism; Key Features: (1) Temporal consistency preservation (2) Multi-frame fusion (3) Motion artifact reduction IoU: 0.714 Private
2 Hao et al. [30] Attention-guided sequential learning framework; Key Features: (1) Channel-wise attention mechanism (2) Frame-to-frame consistency (3) Multi-scale feature integration F1: 0.843 Private
3 Shi et al. [28] UENet: Enhanced GAN-based architecture; Key Features: (1) Adversarial training strategy (2) Multi-level feature enhancement (3) Boundary-aware loss Pixel Acc: 0.8353 Private
4 Xian et al. [59] Deep learning-based main vessel extraction; Key Features: (1) Hierarchical feature learning (2) Smart medical imaging integration (3) Automated vessel tracking F1: 0.900 Private
5 Li et al. [31] CAU-Net with pyramid feature fusion; Key Features: (1) Cascaded attention units (2) Multi-scale pyramid features (3) Deep supervision Dice: 0.904 Private
6 Zhu et al. [60] Enhanced PSPNet architecture; Key Features: (1) Pyramid scene parsing (2) Multi-scale context aggregation (3) Global context modeling Acc: 0.957
Sens: 0.865
Spec: 0.949
Private
7 Thuy et al. [61] Coarse-to-fine strategy for coronary vessel segmentation; Key Features: (1) UNet for the main coronary artery extraction (2) the second UNet for coronary artery-extracted segmentation Dice: 0.7640 Private
8 Tian et al. [62] Hybrid deep learning framework; Key Features: (1) Integrated digital processing (2) Multi-stage refinement (3) Morphological post-processing Dice: 0.939 Private
9 Liu et al. [32] Full-resolution network with dual thresholding; Key Features: (1) High-resolution feature maintenance (2) Adaptive thresholding (3) Multi-dataset validation F1: 0.802 (DCA1)
F1: 0.760 (CHUAC)
DCA1 [49], CHUAC [50]
10 Najarian et al. [63] Ensemble deep learning with filter enhancement; Key Features: (1) Multi-model ensemble strategy (2) Filter-based feature extraction (3) Robust vessel detection F1: 0.874 Private
11 Hau et al. [64] Progressive perception learning framework; Key Features: (1) Stage-wise feature learning (2) Real-time processing (3) Cross-dataset generalization Dice: 0.939 MCS, DCAI [49], XCA [49]
12 Han et al. [29] Recursive centerline-and direction-aware joint learning network; Key Features: (1) centerline- and direction-aware joint learning (2) recursive learning strategy F1: 0.8561 Private
13 Zhang et al. [33] Recursive centerline-aware network; Key Features: (1) Direction-aware feature learning (2) Recursive refinement (3) Ensemble prediction F1: 0.831 Private
14 Navab et al. [34] Weakly-supervised Learning for coronary vessel segmentation Contrastive diffusion-based learning; Key Features: (1) UNet with channel attention module (2) Centerline auxiliary supervision module Dice: 0.8528 Private
15 Zhang et al. [41] Weakly-supervised segmentation with self-paced learning; Key Features: (1) Noisy label handling (2) Curriculum learning (3) Progressive refinement Dice: 0.821 Private
Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union. Key metrics are defined as follows: F1 score is the harmonic mean of precision and recall, ranging from 0 to 1 (max = 1); Dice score (DSC) measures overlap between the predicted output ($A$) and ground truth ($B$) as $DSC=\frac{2|A\cap B|}{|A|+|B|}$, ranging from 0 to 1 (max = 1); AUC is the area under the ROC curve, ranging from 0 to 1 (max = 1); IoU is the intersection over union between the predicted output ($A$) and ground truth ($B$), $IoU=\frac{|A\cap B|}{|A\cup B|}$, ranging from 0 to 1 (max = 1). Additional metrics (AP, Acc, MPE) follow original definitions, with MPE in px or mm.
TABLE III: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
No. Authors Category Method and Key Features Performance Metrics Dataset
16 Qi et al. [65] Examinee-examiner dual network; Key Features: (1) Centerline-guided learning (2) Weak supervision strategy (3) Dual network interaction Dice: 0.762 Private
17 Zhang et al. [66] Semi-supervised cross-anatomy domain adaptation network; Key Features: (1) Vesselness-specific batch normalization (2) Self-ensembling mean teacher Dice: 0.7884 Private
18 Ma et al. [20] Self-supervised adversarial learning; Key Features: (1) Label-free training (2) Adversarial feature learning (3) Consistency regularization Dice: 0.557 XCAD [20]
19 Shen et al. [67] Expert-guided knowledge distillation; Key Features: (1) Semi-supervised learning (2) Expert knowledge transfer (3) Adaptive distillation F1: 0.809 DCA1 [49]
20 Zhang et al. [68] Partial annotation-based segmentation; Key Features: (1) Self-training mechanism (2) Prototype learning (3) Incomplete label handling Dice: 0.715 Private
21 Ma et al. [36] Localize all the endpoints of target vessel segments; Key Features: (1) Coarse-to-fine keypoint detection network (2) Multi-branch feature aggregation (3) Frangi filter with coordinate decoding for boundary enhancement, PCK: 0.763 Private
22 Lee et al. [42] DeNVeR: Deformable neural representations; Key Features: (1) Unsupervised video segmentation (2) Deformable modeling (3) Temporal consistency Dice: 0.733 XCAD [20], CADICA [69], XACV[42]
23 Kim et al. [35] Contrastive diffusion-based learning; Key Features: (1) Label-free diffusion models (2) Contrastive learning (3) Multi-dataset validation Dice: 0.661 (XCAD)
0.700 (134XCA)
0.621 (30XCA)
XCAD [20], 134XCA [49], 30XCA [30]
24 Zhang et al. [70] Curve-like structure detection; Key Features: (1) Efficient Multiscale Module with depthwise separable convolutions (2) triple attention module (3) Boundary enhancement via topological connectivity F1 Score: 0.7912 XCAD[20]
25 Lei et al. [71] Contrastive diffusion-based learning; Key Features: (1) Label-free diffusion models (2) Contrastive learning (3) Multi-dataset validation Dice: 0.661 (XCAD)
0.700 (134XCA)
0.621 (30XCA)
XCAD [20], CHUAC [50], DCA1 [49]
26 Iyer et al. [21] Clinical Applications Systematic anatomical assessment of coronary stenosis in the clinical workflow; Key Features: (1) Deeplabv3+ (2)Clinical validation Dice: 0.864 Private
27 Song et al. [43] Width-aware UNet for clinical assessment; Key Features: (1) Vessel width prediction (2) Enhanced UNet architecture (3) Clinical validation Dice: 0.773 Private
28 Yao et al. [25] Integrated path planning system; Key Features: (1) Heuristic path optimization (2) Clinical workflow integration (3) Real-time guidance F1: 0.755 Private
29 Shen et al. [72] Vascular segmentation for clinical use; Key Features: (1) UNet with bidirectional ConvLSTM (2) Clinical workflow integration F1: 0.879 Private
30 Nobre et al. [44] Multi-center clinical validation study; Key Features: (1) Large-scale validation (2) Multi-center data collection (3) Clinical outcome assessment Dice: 0.948 Private
31 Wang et al. [46] CE-NC-VesselSegNet for clinical use; Key Features: (1) Cross-modality learning (2) Transfer learning strategy (3) Clinical cohort validation Dice: 0.856 (cohort 1)
0.738 (cohort 2)
Private
32 Yang et al. [47] Geometry-constrained cascaded network; Key Features: (1) Geometric prior integration (2) Multi-stage refinement (3) Cross-dataset validation Dice: 0.778 (ASOCA)
0.895 (CCA-200)
ASOCA [48], CCA-200
Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
TABLE IV: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
No. Authors Category Method and Key Features Performance Metrics Dataset
33 Fu et al. [73] Robust temporal foreground extraction; Key Features: (1) Temporal consistency modeling (2) Noise-resistant features (3) Adaptive thresholding F1: 0.801 Private
34 Guo et al. [45] Edge-preserving vessel segmentation; Key Features: (1) Scribble-based supervision (2) Edge enhancement (3) Boundary refinement Dice: 0.734 XCAD [20]
35 Lalinia et al. [74] Edge-guided tracking framework; Key Features: (1) Multi-scale edge detection (2) Tracking optimization (3) Clinical integration Dice: 0.818 Private
36 Zhou et al. [75] Guidewire Detection and Tracking Pyramid attention recurrent network; Key Features: (1) Real-time processing (2) Multi-scale attention (3) Temporal modeling Dice: 0.942 Private
37 Zhou et al. [22] Lightweight recurrent attention framework; Key Features: (1) Resource-efficient design (2) Real-time capability (3) Attention mechanism Dice: 0.947 Private
38 Zhou et al. [76] two stage multifunctional framework for guidewire endpoint localization; Key Features: (1) fast attention recurrent network (2) endpoint localization and angle measurement F1 Score: 0.938, mean distance error: 0.596 pixels Private
39 Li et al. [77] Multi-guidewire endpoint detection; Key Features: (1) Multiple wire tracking (2) Endpoint localization (3) Unified detection framework AP: 0.9065 Private
40 Li et al. [78] Gaussian-based probabilistic modeling; Key Features: (1) Probabilistic representation (2) Teleoperation guidance (3) Uncertainty estimation Acc: 0.998 Private
41 Zhu et al. [79] Multi-stage tip tracking system; Key Features: (1) Cascaded processing (2) Precise tip localization (3) Real-time tracking F1: 0.915 Private
42 Zhang et al. [37] Real-time intraoperative tracking; Key Features: (1) Online tracking (2) Low latency (3) Clinical workflow integration Dice: 0.899 Private
43 Zhang et al. [80] Jigsaw transformer framework; Key Features: (1) Background reverse attention (2) Jigsaw training strategy (3) Multi-wire handling F1: 0.899 (single)
0.846 (dual)
Private
44 Ravigopal et al. [81] Real-time pose estimation system; Key Features: (1) Continuum kinematic modeling (2) Real-time tracking (3) Pose optimization Acc: 0.900 Private
45 Wen et al. [82] Sim-to-real transfer framework; Key Features: (1) Domain adaptation (2) Synthetic data utilization (3) Cross-domain generalization IoU: 0.660 Cardiac [38]
46 Du et al. [16] Guidewire endpoint localization via pixel-adjacent relation; Key Features: (1) UNet segmentation (2) Eight-neighborhood search mean pixel errors: 2.02 Private
47 Gherardini et al. [38] Catheter Detection and Tracking Lightweight UNet with synthetic data; Key Features: (1) Transfer learning (2) Efficient architecture (3) Synthetic data augmentation Dice: 0.875 Cardiac [38]
48 Kolen et al. [83] DQN-driven segmentation system; Key Features: (1) Deep Q-learning (2) Dual UNet architecture (3) Self-supervised training Dice: 0.650 Private
49 Zhou et al. [39] Fast recurrent residual network; Key Features: (1) Recurrent processing (2) Residual connections (3) High-speed inference F1: 0.950 Private
50 Omisore et al. [84] Robotic tool segmentation; Key Features: (1) Robot-assisted tracking (2) Real-time processing (3) Interventional integration AUC: 0.753 Private
Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
TABLE V: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
No. Authors Category Method and Key Features Performance Metrics Dataset
51 Aghasizade et al. [40] Coordinate regression detection; Key Features: (1) Precise localization (2) Structural intervention (3) Coordinate prediction Mean Error: 7.36 px Private
52 Ranne et al. [85] CathFlow: Self-supervised framework; Key Features: (1) Optical flow guidance (2) Transformer architecture (3) Unsupervised learning Dice: 0.728 Private
53 Ma et al. [86] Dynamic roadmapping system; Key Features: (1) Bayesian filtering (2) Deep learning fusion (3) Real-time mapping Mean Error: 0.96 mm Private
54 Wu et al. [87] multi-modal Registration CAR-Net: 3D/2D registration framework; Key Features: (1) Deformation modeling (2) Real-time registration (3) Multi-view fusion Mean Error: 1.13 mm Private
55 Yan et al. [88] Deep reinforcement learning-based multi-modal registration; Key Features: (1) deep reinforcement learning (2) reward function: euclidean distance MPE: 2.6544 (LCA), 2.4825 (RCA) Private
56 Jaganathan et al. [89] Self-supervised multi-modal registration; Key Features: (1) simulated training with unsupervised feature (2) pixel space domain adaptation Mean Error: 1.83 $\pm$ 1.16 mm Private
57 Huang et al. [90] 2D/3D Registration via CNN Regression and Centroid Alignment; Key Features: (1) CNN-based transformation parameters estimation (2) Vessel centroids refinement Dice: 0.7669, 0.785, and 0.8539 on three patients respectively Private
58 Liu et al. [91] Deep reinforcement learning-based rigid registration; Key Features: (1) overlap degree calculation reward function (2) optimizing spatial transformation parameters via online learning MAE: 2.85 mm Private
Note: AP = Average Precision; Acc = Accuracy; AUC = Area Under Curve; px = pixels; mm = millimeters; MPE = Mean Projective Error. Performance metrics are reported as in the original papers. Single/dual refers to single or dual guidewire scenarios.
II-B1 Guidewire Detection and Tracking
Guidewire detection and tracking are essential components in minimally invasive endovascular procedures, facilitating precise navigation through vascular pathways prior to catheter insertion. As the initial interventional tool to be navigated, the guidewire ensures safe access to the target site and establishes a stable platform for subsequent catheter placement. Recent advancements in this field have concentrated on enhancing the accuracy and real-time performance of surgical instruments tracking, ultimately improving its clinical applicability[92, 93, 94], through the integration of deep learning techniques. These innovations increase guidance precision, facilitating effective navigation through tortuous vasculature and contributing to safer, more efficient procedures [81].
Zhou et al. [75] introduced a pyramid attention recurrent network that employs multi-scale attention mechanisms and temporal modeling to achieve real-time guidewire detection. Their method exhibited impressive performance levels, achieving a Dice coefficient of 0.942 on a private dataset. Building on this research, Zhou et al. [22] proposed a lightweight recurrent attention framework that emphasizes resource efficiency while maintaining high accuracy (Dice: 0.947). This approach provides a feasible solution for optimizing the critical trade-off between performance and computational efficiency.
Addressing the challenge of multi-guidewire scenarios, Li et al. [77] developed a unified detection framework capable of tracking multiple guidewires and localizing their endpoints. Their method achieved an average precision of 0.9065 on a private dataset, marking significant progress in complex interventional procedures. In a subsequent study, Li et al. [78] introduced a Gaussian-based probabilistic modeling approach, which not only provided highly accurate tracking (Accuracy: 0.998) but also incorporated uncertainty estimation, thereby enhancing the system’s reliability in relation to tele-operative guidance.
Zhu et al. [79] focused on precise tip localization using their multi-stage tracking system, achieving an F1 score of 0.915. This cascaded processing approach demonstrates the potential for highly accurate real-time tracking of guidewire tips, a crucial aspect in numerous interventional procedures.
Recent research has emphasized clinical integration and real-time performance. Zhang et al. [37] developed a system for real-time intra-operative tracking with low latency (Dice: 0.899), addressing the critical need for seamless integration into clinical workflows. Their subsequent work [80] introduced the Jigsaw transformer framework, incorporating background reverse attention and a novel training strategy to effectively handle multi-wire scenarios (F1: 0.899 for single wire, 0.846 for dual wire).
Ravigopal et al. [81] advanced the field with a real-time pose estimation system that combines continuum kinematic modeling with pose optimization, achieving an accuracy of 0.900. This approach underscores the potential for more comprehensive guidewire tracking beyond simple detection.
Addressing the limitations of insufficient clinical data, Wen et al. [82] proposed a sim-to-real transfer framework that utilizes synthetic data and domain adaptation techniques to improve generalization. Their method achieved an IoU of 0.660 on the public Cardiac dataset [38], demonstrating the potential of leveraging synthetic data to enhance guidewire detection models.
II-B2 Catheter Detection and Tracking
Catheter detection and tracking constitute critical components in minimally invasive cardiovascular procedures, directly enhancing procedural accuracy, facilitating real-time guidance, and minimizing the risk of complications [94, 95]. These technologies enable more precise navigation through complex vascular networks by providing continuous and accurate tracking of the catheter’s position and movement [19, 96], and therefore reducing the likelihood of inadvertent damage to surrounding tissue [97]. Furthermore, they facilitate better targeting of therapeutic areas, such as lesions or thrombi, ultimately leading to enhanced treatment outcomes and shortened procedure times. Recent innovations in this field have led to significant advancements in both algorithmic accuracy and practical applications, offering tailored solutions to address the unique challenges of cardiovascular interventions.
Gherardini et al. [38] proposed a lightweight UNet architecture that leverages synthetic data and transfer learning for catheter detection in medical images. Their approach achieved a Dice coefficient of 0.875 on the Cardiac dataset, demonstrating the effectiveness of synthetic data augmentation and efficient network design. Yang et al. [83] introduced a semi-supervised deep Q-network-driven segmentation method for 3D ultrasound images, enabling accurate real-time catheter detection and tracking. This novel method, which combines deep Q-learning with a dual UNet architecture, obtained a Dice score of 0.650 on a private dataset, demonstrating the feasibility of reinforcement learning in medical image analysis. Zhou et al. [39] presented FRR-Net, a fast recurrent residual learning-based architecture for catheter tracking. This method achieved an impressive F1 score of 0.950 on a private dataset, highlighting the benefits of combining recurrent processing with residual connections for high-speed inference. Collectively, these approaches demonstrate the potential of deep learning models to accurately detect catheters while minimizing computational requirements.
Omisore et al. [84] developed an automatic tool segmentation method that integrates robotic systems and real-time tracking, enabling precise catheter localization during interventions. Their system achieved an AUC of 0.753 on a private dataset, emphasizing the potential of robot-assisted tracking in endovascular procedures. Aghasizade et al. [40] proposed a coordinate regression detection approach that enables precise localization and structural intervention, achieving a mean error of 7.36 pixels on their private dataset. Ranne et al. [85] introduced CathFlow, a self-supervised segmentation method utilizing optical flow and transformers for accurate catheter detection and tracking. This innovative approach, leveraging unsupervised learning techniques, achieved a Dice score of 0.728 on a private dataset, demonstrating the potential of self-supervised learning in medical image analysis. Ma et al. [86] developed a dynamic road mapping system that combines Bayesian filtering with deep learning fusion for real-time catheter mapping. Their method achieved a mean error of 0.96 mm, showcasing its potential for highly accurate catheter localization in clinical settings.
II-C Multi-modal Registration and Integration in Cardiovascular Imaging
Multi-modal registration and integration are crucial in cardiovascular imaging, as they provide enhanced accuracy in diagnosis and treatment [98]. The accurate alignment of high-resolution, real-time DSA with detailed three-dimensional (3D) anatomical information from CTA can facilitate a more comprehensive understanding of complex cardiac structures [23, 87]. The combination of these modalities improves spatial alignment, thereby enhancing vascular navigation and visualization, ultimately leading to more effective procedures and better clinical outcomes [26].
Dynamic coronary road mapping has significant clinical implications, particularly in enhancing the precision of coronary interventions. Ma et al. introduced a novel approach leveraging Bayesian filtering and deep learning fusion [86]. This method dynamically integrates multi-modal imaging data, providing real-time guidance during coronary procedures. The Bayesian filtering technique effectively manages the uncertainty and noise inherent in dynamic imaging, while deep learning fusion integrates various modalities to produce a coherent and precise roadmap. This integration is particularly beneficial in overcoming the limitations of traditional non-learning-based methods, which often struggle with real-time adaptability and accuracy.
Wu et al. presented CAR-Net, an advanced system for 3D/2D registration that is crucial for aligning preoperative 3D images with intra-operative 2D images [87]. This technique is essential for accurate navigation during minimally invasive endovascular procedures. CAR-Net employs sophisticated deformation modeling alongside real-time capabilities, making it a robust tool in clinical settings. Deformation modeling ensures that registration accounts for anatomical changes and movements occurring during the procedure - challenges that traditional methods often fail to address effectively. The real-time capability of CAR-Net further enhances its usability, providing immediate feedback and adjustments during operation. Such learning-based methods surpass traditional approaches by offering higher accuracy and adaptability, which are critical in dynamic and complex clinical environments.
TABLE VI: Summary of Data-Driven Control Methods for Endovascular Intervention Robots
No. Authors (Year) Task Category Method Performance Metrics
1 Chi et al. (2020) [99] Catheter Navigation and Control GAN-based imitation learning: - WGAN architecture - Collaborative robot control - Real-time trajectory generation - Success rate - Learning efficiency - Real-time performance
2 Kweon et al. (2021) [100] Guidewire Navigation (Static Environment) DRL framework: - PPO algorithm - CNN-based state encoder - Continuous action space - Task completion rate - Generalization ability - Path accuracy
3 Jolaei et al. (2021) [101] Catheter Control (Cardiac Ablation) Learning-based kinematics: - Neural network controller - Tendon-driven model - Real-time adaptation - Control accuracy - Position error - Response time
4 Karstensen et al. (2022) [102] Guidewire Navigation (Venous System) Autonomous navigation system: - Vision-based tracking - Path planning algorithm - Real-time control - Success rate - Navigation precision - Safety metrics
5 Yang et al. (2022) [103] Guidewire Control (Feeding Task) DRL-based control: - DDPG algorithm - Custom reward design - State-action mapping - Task success rate - Control efficiency - Stability
6 Zhou et al. (2022) [104] Interventional Skill Learning Skill learning framework: - Feature extraction - Behavior cloning - Skill transfer - Learning efficiency - Task accuracy - Skill transfer rate
7 Dreyfus et al. (2022) [105] Magnetic Guidewire Navigation RL-based navigation: - Magnetic field control - Path optimization - Safety constraints - Position error - Navigation time - Safety index
8 Li et al. (2023) [106] Interventional Skill Learning Conservative actor-critic: - Smooth gradient policy - Stability optimization - Experience replay - Learning stability - Task success rate - Convergence speed
9 Tian et al. (2023) [107] Catheter Navigation (Virtual Environment) DDPG-based method: - Virtual training environment - Continuous control - Transfer learning - Navigation accuracy - Real-time performance - Generalization
11 Li et al. (2024) [108] Guidewire Navigation (Autonomous Delivery) Model-based offline RL: - Conservative Q-learning - Dynamics modeling - Safety constraints - Completion time - Safety metrics - Control stability
12 Mei et al. (2024) [109] Guidewire Navigation (Autonomous Delivery) Actor-critic architecture - Hierarchical learning - Knowledge transfer - Success rate - Learning speed - Decision accuracy
13 Wang et al. (2024) [110] Guidewire Navigation (Autonomous Control) Deep RL framework: - Custom network architecture - Reward shaping - Safety-aware policy - Task success - Learning efficiency - Safety index
14 Jianu et al. (2024) [111] Simulation and Training Open-source simulator: - Physics-based modeling - Real-time simulation - Training environment - Simulation fidelity - Real-time execution - Training effectiveness
15 Li et al. (2024) [108] Vascular Navigation (Autonomous) Discrete SAC with AE: - Auto-encoder state space - Discrete action mapping - Policy optimization - Learning efficiency - Task success rate - Robustness
16 Robertshaw et al. (2024) [112] Intra-vascular Navigation DRL-based navigation: - End-to-end learning - Vision-based control - Safety constraints - Task success rate - Navigation efficiency - Safety metrics
TABLE VII: Summary of Data-Driven Control Methods for Endovascular Intervention Robots (continued)
No. Authors (Year) Task Category Method Performance Metrics
17 Scarponi et al. (2024) [113] Guidewire Navigation (Dynamic Environment) Autonomous navigation: - Dynamic environment handling - Real-time adaptation - Robust control - Navigation accuracy - Robustness - Adaptability
18 Liu et al. (2024) [114] Image-guided Navigation Vision-based RL: - Image processing pipeline - Multi-modal fusion - Real-time control - Navigation accuracy - Precision - Real-time performance
19 Scarponi et al. (2024) [115] Guidewire Navigation (Zero-shot Learning) Zero-shot RL strategy: - Domain adaptation - Policy transfer - Dynamic environment handling - Navigation accuracy - Adaptation ability - Robustness
20 Yao et al. (2025) [116] Guidewire Navigation (Simulation to reality) Autonomous navigation system: - Domain adaptation - Policy transfer - Dynamic environment handling - Navigation accuracy - Task accuracy - Adaptability
Figure 6: Learning frameworks for robotic-assisted endovascular procedures can be categorized as follows: (a) The traditional reinforcement learning (RL) framework for endovascular navigation involves a robotic agent learning through direct interaction with the vascular environment. This approach utilizes target networks and experience buffers to optimize catheter/guidewire manipulation policies [4, 102, 96, 107, 115]. (b) The demonstration-based learning framework incorporates expert cardiologist demonstrations to guide the learning of robotic manipulation policies. This method enables more efficient acquisition of complex endovascular navigation skills [106, 104, 108]. (c) The human-in-the-loop learning framework integrates expert cardiologist supervision into the robotic training process. This approach combines automated learning with clinical expertise to enhance safety and efficacy in endovascular procedures [111, 117].
Figure 7: Demonstration of Learning-based control methods for endovascular interventional navigation. (a) Virtual training for guidewire navigation [102]. (b) Autonomous guidewire navigation using inverse reinforcement learning [112]. (c) Virtual environment based on 2D DSA images [109]. (d) Control strategy for deforming vascular anatomy [113]. (e) Catheter insertion interventional simulator based on finite element simulation [117]. (f) Interventional operation using joystick control data collection [106]. Images reproduced from [102], [112], [109], [113], [117], and [106].
Building upon these developments, several innovative approaches have emerged to address specific challenges in medical image registration. Jaganathan et al. [89] introduced a self-supervised multi-modal registration framework that leverages unsupervised feature learning and pixel-space domain adaptation. This approach demonstrated remarkable precision, achieving a mean error of 1.83 ± 1.16 mm, thereby highlighting the potential use of self-supervised learning in handling complex registration tasks without the need for extensive manual annotations. Alongside this, Huang et al. [90] developed a sophisticated 2D/3D registration method that combines convolutional neural network (CNN) regression with centroid alignment. Their dual-stage approach employs CNN-based transformation parameter estimation, followed by precise vessel centroid refinement. This methodology achieved impressive Dice scores ranging from 0.7669 to 0.8539 across different patient cases, demonstrating robust performance in clinical applications. Furthermore, Liu et al. [91] proposed an innovative deep reinforcement learning framework for rigid registration. Their method introduces a novel reward function based on overlap degree calculation and optimizes spatial transformation parameters through online learning. With a mean absolute error of 2.85 mm, this approach exemplifies how reinforcement learning can effectively navigate the complex space of registration parameters while maintaining clinical accuracy.
III Learning-Based Control for Robotic-Assisted Endovascular Procedures
Endovascular devices constitute a unique class of continuum robots, characterized by their inherent flexibility and adaptability in navigating complex vascular networks [118]. The fundamental challenges posed by these devices arise from three key factors: the nonlinear dynamics of flexible instruments [119, 120], the variability of vascular anatomy [2], and the limited real-time visualization during procedures. Traditional rigid-body control methods prove inadequate in addressing these challenges, particularly when precise manipulation is required within confined, tortuous pathways [121, 122].
Learning-based control strategies have emerged as promising solutions, offering several advantages over conventional approaches [123]. These methods can effectively capture the complex relationships between device inputs and outputs [124, 51], adapt to patient-specific anatomical variations, and optimize real-time decision-making during interventions. By incorporating data-driven techniques, these control systems can enhance procedural precision while reducing the cognitive load on cardiologists, ultimately improving the safety and efficiency of endovascular procedures [125].
The development of control systems has traditionally relied heavily on real-world data collection and validation. Early studies emphasized task-specific methodologies, leveraging real-world experimental setups to refine algorithms and demonstrate clinical applicability. For instance, Kweon et al. [100] established benchmarks in coronary artery navigation through their deep reinforcement learning (DRL) framework, achieving high task completion rates and path accuracy. Similarly, Zhou et al. [104] developed intervention skill learning frameworks based on feature extraction and behavior cloning, capturing expert manipulation characteristics with remarkable efficiency. Despite these successes, real-world data collection remains limited by practical challenges, such as patient variability [126], ethical concerns, and the high cost of clinical experimentation.
To overcome these limitations, researchers have increasingly turned to virtual environments, which offer significant advantages in algorithm training and validation. Simulated platforms enable safe, repeatable, and cost-effective experimentation, circumventing the challenges associated with real-world data collection. In one such example Jianu et al. [111] introduced CathSim, an open-source simulator providing high-fidelity physics-based modeling for real-time endovascular procedural training. Similarly, Li et al. [106] employed a conservative actor-critic method in a simulation environment, achieving high learning stability and task success rates even in complex endovascular procedure scenarios. Yang et al. [103] further demonstrated the effectiveness of simulation-based training with their DRL control framework, optimizing control efficiency in guidewire feeding tasks.
The value of virtual environments is further underscored by advancements in other areas of robotics. Quadruped and humanoid robots have, for example, utilized the NVIDIA OMNIVERSE platform to conduct skill learning in highly realistic and diverse simulated environments, enabling scalable training across various tasks. In surgical robotics, the Da Vinci system has leveraged platforms like SurRoL [127] and Surgical Gym [128] for virtual training, effectively replicating intricate endovascular procedures. These examples illustrate how simulation platforms can accelerate algorithm development and reduce the dependence on scarce real-world data.
Building on these advantages, the emerging trend of Simulation to Reality (Sim2Real) has become a critical focus in endovascular procedural research [129]. Sim2Real methodologies aim to bridge the gap between simulation and real-world applications, ensuring that algorithms trained in virtual environments generalize effectively to clinical scenarios. Recent studies highlight significant progress in this area. Tian et al. [107] employed domain randomization techniques in a Deep Deterministic Policy Gradient (DDPG)-based framework, enhancing generalization and real-time performance. Scarponi et al. [115, 113] further advanced zero-shot learning methods, demonstrating robust adaptability in dynamic environments.
The integration of simulation platforms with real-world data has also enabled hybrid approaches that combine the strengths of both methodologies. Wang et al. [110] and Liu et al. [114] developed multi-modal image-guided navigation systems, achieving enhanced precision and adaptability under clinical conditions. Dreyfus et al. [105] applied reinforcement learning (RL) to magnetic guidewire navigation, incorporating safety constraints and path optimization strategies.
As research in the field continues to progress, virtual environments are expected to play an increasingly central role in training autonomous systems. By reducing the reliance on real-world data and enabling efficient algorithm development, simulation platforms, coupled with Sim2Real advancements, are poised to transform the landscape of robotic-assisted endovascular procedures.
IV Challenges
Data-driven approaches in endovascular robotic systems offer considerable promise, yet several critical challenges persist that need to be addressed to fully harness their potential. These challenges lie within the technical, ethical, and systemic domains - areas in which substantial progress is necessary to advance key capabilities and clinical adoption.
IV-A Underdevelopment of Transformer-Based Models and Foundational Vision Models for Endovascular Procedures
Despite significant advancements in the use of transformer-based architectures and large pre-trained vision models in medical imaging and robotics [130, 131], their integration into endovascular procedures remains largely unexplored. Transformers, which have achieved remarkable success in domains such as natural language processing and computer vision [132], offer substantial advantages when capturing long-range dependencies and processing complex, high-dimensional data. These characteristics render transformers particularly well-suited to dynamic visual tasks in endovascular procedures, including real-time vascular segmentation, lesion detection, and procedural workflow recognition. However, their full potential in this specialized field has yet to be fully realized. A key reason for this is the unique challenges posed by endovascular data. Unlike other imaging tasks, endovascular images are often noisy and exhibit significant inter-patient variation [133, 134], compounded by complex and highly variable topology [25, 17, 135]. These factors necessitate models that can not only accurately delineate intricate anatomical structures but also handle the temporal dynamics of procedures, such as real-time vessel navigation, tool positioning, and tissue response. Despite the demonstrable success of large-scale transformer-based models in general AI tasks, their application in medical vision tasks requiring high spatiotemporal accuracy, particularly in the specialized domain of coronary and endovascular imaging, remains insufficiently explored [136]. Alongside this, the integration of transformer models with multi-modal inputs, such as the combination of preoperative imaging with intra-operative fluoroscopy or ultrasound, is also an area that has seen limited development in relation to endovascular research.
A significant barrier to the development of transformer-based solutions for endovascular procedures is the scarcity of large-scale, high-quality annotated datasets necessary for training such models [137]. While transformers excel when trained on vast amounts of data, current endovascular imaging datasets are often small, highly specialized, and lack sufficient diversity in patient populations, anatomical variations, and procedural conditions. This limits the utilization of transformer models in this domain. Furthermore, the transferability of large pre-trained models from other domains, such as chest X-rays [138], brain MRIs [139], or general robotic manipulation [140], into endovascular imaging remains a challenge. To address this, future research should prioritize the creation of large, multi-institutional, and multi-modal datasets that capture the broad diversity of patient anatomies, procedural contexts, and pathological conditions, all of which are crucial for training robust and generalizable models.
IV-B Scarcity of High-Quality Coronary Imaging Data
Despite the extreme importance of having high-quality reference data in coronary imaging, current literature reveals surprisingly little research on this. The inherent complexity of coronary arteries, combined with significant variability in patient presentation and imaging conditions, necessitates robust gold standards and extensive, well-curated datasets for reliable AI development. However, the acquisition of such data is problematic, and certain hurdles need to be overcome. In the first place, coronary imaging data often requires specialized expertise for annotation. Moreover, the dynamic nature of coronary arteries, influenced by cardiac motion and respiratory movements, adds another layer of complexity to data acquisition and annotation. While technological advancements in AI algorithms continue to emerge, empirical evidence suggests that in the intricate domain of coronary imaging, data quality is often more crucial than algorithmic sophistication in determining model performance [136, 141]. Essentially, we need coordinated efforts to establish standardized protocols for coronary imaging data collection, annotation, and sharing across institutions. This challenge is further compounded by the necessity to capture diverse pathological presentations, varying degrees of stenosis, different plaque characteristics, and the full spectrum of anatomical variations in coronary arteries. Addressing these data-related challenges is fundamental to advancing AI applications in coronary imaging and, ultimately, improving cardiovascular care through AI.
IV-C Lack of Unified Workflow Recognition for Endovascular Procedures
The absence of a unified approach to workflow recognition in endovascular procedures poses a critical challenge, significantly impeding the advancement of robotic-assisted procedures and the development of higher-level autonomy in medical robotics [142, 143, 144, 145]. In contrast to laparoscopic surgery, where initiatives such as Autolaparo [146] and CholecTriplet2022 [147] have provided comprehensive datasets and benchmarks, endovascular procedures lack comparable resources. This deficiency critically inhibits progress towards next-generation intervention systems capable of autonomous navigation, adaptive force control, automated decision-making, and intelligent task planning.
Accurate and real-time workflow recognition is paramount in robotic-assisted endovascular procedures, particularly in managing the complex sequence of instrument usage and ensuring the proper selection and timely exchange of tools, such as guidewires, catheters, and stents, throughout various procedural phases. By anticipating transitions between interventional steps, it facilitates seamless instrument switches, reduces procedure time, and enhances overall efficiency and safety in complex procedures. This technology enables robotic systems to proactively prepare and present the correct instrumentation, thereby minimizing delays and mitigating thrombotic complications associated with prolonged intervention. In emergency scenarios requiring rapid changes in instrumentation, such as deploying a rescue stent during vessel dissection, swift phase recognition can trigger immediate tool preparation and exchange [148, 142]. By augmenting a robot’s ability to anticipate and adapt to evolving instrument needs across procedural stages, this technology substantially improves both efficiency and safety, representing a pivotal step toward achieving intelligent endovascular robotic systems capable of high-level decision-making and adaptive behavior in dynamic environments.
IV-D Absence of a Standardized, Large-Scale Benchmark
Recent advancements in large-scale benchmarking, exemplified by the Behavior-1K dataset [149] developed by Fei-Fei Li’s team, have significantly enhanced performance evaluation in endovascular interventions (EIs) by capturing 1,000 diverse, real-world activities within controlled simulation environments. The work deemed "Best Paper at ICRA 2024" also extended robotic learning through the introduction of standardized, high-quality datasets and models that support reproducibility and generalization across a wide range of robotic systems [150]. In contrast, endovascular robotics remains constrained by the scarcity of large-scale, domain-specific benchmarks that encapsulate the intricacies of human vasculature and the clinical complexity encountered in real-world settings. This limitation hampers systematic comparisons of algorithmic performance, impedes meaningful evaluation under varying clinical conditions, and results in fragmented, non-generalizable findings.
Although emerging efforts, such as CathAction [151] and stEVE [152], aim to advance benchmark frameworks for understanding and learning in endovascular procedures, the field still lacks datasets that reflect the breadth of real-world, multi-modal, and multitask scenarios commonly encountered in clinical practice [153]. The absence of a comprehensive, standardized evaluation environment fragments the development and validation of endovascular control algorithms, thereby hindering scientific innovation and slowing the translation of advanced solutions into clinical utility.
Addressing this critical gap necessitates the development of a benchmarking framework specifically tailored to endovascular robotics, incorporating diverse anatomical models, robust performance metrics, and high-fidelity simulations that accurately represent physiological dynamics. Such an infrastructure would facilitate consistent, quantitative assessments of algorithmic performance across various measures, including trajectory precision, navigation efficiency, collision avoidance, and procedural safety. Drawing inspiration from established benchmarking efforts in robotic manipulation and EI could provide valuable methodological guidance, ultimately fostering a cycle of iterative improvement. By taking these steps, one could enhance both the technical sophistication, and clinical relevance, of endovascular robotic control, paving the way for better, safer, seamlessly integrated clinical procedures.
IV-E Scalability of Perception and Learning-Based Control
Scalability remains a pivotal challenge for perception systems in endovascular procedures, given that these systems need to navigate the complex and dynamic vascular environment. In contrast to traditional robotic domains, endovascular robotics operates within an extremely intricate and continuously evolving anatomical landscape, characterized by varying vessel sizes, tortuous paths, and diverse patient-specific conditions. These factors substantially complicate the challenge of realizing accurate and reliable real-time perception. Furthermore, unlike rigid environments, the vascular system undergoes significant physiological change during procedures, including vessel deformation and blood flow fluctuations. Consequently, existing perception systems often struggle to maintain consistent accuracy and robustness across diverse patient anatomies, thereby increasing the risk of procedural errors.
This highly variable and frequently unpredictable nature of patient anatomy and pathology poses particular adaptability-related challenges for learning-based control systems. These typically rely on data-driven models trained on a limited subset of patient cases, impeding their ability to generalize effectively across the broad spectrum of anatomical variations and disease states encountered in clinical practice. Systems trained on the anatomy of one patient may fail to perform optimally when applied to a new patient with a differing vascular anatomy. This lack of generalizability significantly reduces their effectiveness in real-world environments.
To address these scalability challenges, future research must concentrate on developing advanced perception algorithms capable of maintaining high accuracy and robustness across a wider range of anatomical variations and pathological conditions. One promising direction involves the integration of multi-modal imaging techniques, including ultrasound and DSA [154], combined with advanced deep learning algorithms, to enhance the system’s capacity to adapt to real-time physiological changes. Additionally, transfer learning and domain adaptation approaches may aid in overcoming the challenge of generalizing control algorithms across different patient datasets. These strategies have the potential to substantially improve the flexibility and clinical applicability of endovascular robotic systems, ensuring their reliable performance across diverse patient populations, and ultimately leading to improved procedural outcomes and patient safety.
IV-F Ethical and Legal Concerns
The integration of autonomous robotic systems into medical procedures poses several ethical and legal challenges that must be addressed prior to their implementation. As robotics plays an increasing role in high-stakes procedures such as endovascular interventions, it is imperative to consider key issues surrounding liability, decision-making autonomy, patient consent, and data privacy, alongside the principles of trustworthiness and explainable artificial intelligence (XAI) [155].
Liability is a paramount concern, given the possibility of autonomous robots making clinical decisions that result in errors [156]. In traditional medical practice, responsibility is assigned to healthcare providers. However, in the context of autonomous systems, determining accountability becomes complex. Clear legal frameworks are necessary to delineate liability and ensure that the relevant stakeholders are held accountable in the event of adverse outcomes.
Decision-making autonomy represents another critical issue. While AI-powered robotic systems can enhance precision, their "closed-box" nature, with opaque decision-making processes, raises concerns about clinician trust [157, 158]. Trustworthiness, defined as the reliability and predictability of a system, is crucial to ensure clinicians have confidence in the technology’s performance. Furthermore, XAI is essential for transparency, enabling clinicians to understand the rationale behind the robot’s decisions. This is particularly important in high-risk procedures, in which clinicians may need to intervene if the system’s decision-making process is unclear [159].
Patient consent becomes more nuanced when robotic systems are involved [160]. Patients must be informed not only about the risks and benefits of a procedure but also about the role of AI in decision-making. Clear and comprehensible explanations are essential to ensure patients understand precisely how autonomous systems will be involved in their care.
Data privacy also poses a significant challenge in the development of autonomous medical systems, as these technologies rely heavily on extensive datasets for training and continuous algorithmic refinement [160]. Ensuring robust protection and ethical utilization of patient data is of utmost importance, extending beyond technical considerations to encompass fundamental human rights and medical ethics [161]. Comprehensive ethical guidelines must be carefully constructed to govern the use of patient data, with a primary focus on ensuring complete transparency and patient autonomy over the use of any sensitive information for algorithmic training. These guidelines should provide clear, granular mechanisms for informed consent, data anonymization, and stringent access controls. Federated learning emerges as a promising technological solution to this complex privacy dilemma, offering a sophisticated approach that enables collaborative model training while preserving the confidentiality of individual patient records [162]. By allowing algorithms to learn from distributed datasets without centralizing raw patient information, federated learning represents an innovative paradigm that can potentially reconcile the seemingly conflicting objectives of advanced medical AI and patient privacy protection.
The issue of equity also needs to be addressed. The high cost of robotic systems may limit their accessibility, potentially exacerbating disparities in healthcare provision. Efforts should be made to ensure equitable access to these technologies, ensuring they benefit a diverse range of patient populations.
IV-G Integration with Existing Clinical Workflows
Beyond the technical challenges, the integration of robotic systems into existing clinical workflows poses a critical barrier to their widespread adoption. Although the potential of endovascular robotics to enhance precision, decrease procedural variability, and improve patient outcomes is evident, the transition from traditional manual methods to fully automated or semi-automated systems presents substantial challenges beyond purely technological aspects. A pivotal issue is the reluctance of clinical staff to embrace new systems, particularly when they are perceived as "black boxes" - autonomous systems with decision-making processes that are difficult to interpret or explain [158]. In endovascular procedures, in which precision and safety are paramount, the inability to fully comprehend how a robotic system arrives at its decisions can elicit significant levels of anxiety among medical professionals. Clinicians require assurances that these systems can manage complex and unpredictable situations in a manner akin to experienced human practitioners, and that they would operate predictably in high-pressure clinical environments [163].
Another challenge involves the interoperability of robotic systems with existing medical infrastructure and established clinical workflows [164]. Endovascular procedures rely on a sophisticated network of medical technologies, including imaging systems (e.g., fluoroscopy, ultrasound, and CT), patient monitoring systems, and intervention tools, which all function seamlessly together. For successful integration, robotic systems must offer interoperability with these platforms, facilitating efficient and reliable communication without disrupting the workflow or necessitating substantial retraining of medical staff. Standardization of data formats, communication protocols, and system interfaces is crucial to ensure that these systems can be readily adopted and maintained within a clinical setting.
Furthermore, the integration process should consider not only technological compatibility but also workflow alignment. Robotic systems should be designed to complement existing clinical practices [165], minimizing disruption to the time-sensitive nature of endovascular procedures. This includes ensuring that they can be swiftly and easily incorporated into procedure setups, do not excessively complicate or prolong interventions, and provide real-time feedback in a format accessible to clinicians in the operating room. Ultimately, automated guidance and decision support systems should augment, rather than supplant, physician expertise, facilitating enhanced collaboration between human and robotic components during procedures.
Training and education are essential factors in successful integration. Medical professionals not only need to acquire proficiency in operating these systems but also need to develop confidence in their use, along with a clear understanding of their limitations [166, 167]. Training programs should emphasize both technical skills and clinical decision-making processes, empowering healthcare providers when working alongside robotic systems. Without this educational investment, even the most advanced robotic technologies risk under-utilization in clinical practice.
V Conclusion
The integration of data-driven methodologies into endovascular robotics signifies a paradigm shift in minimally invasive endovascular procedures. This comprehensive review systematically analyzes the convergence of advanced robotics, artificial intelligence (AI) algorithms, and multi-modal medical imaging technologies, demonstrating their collective potential to augment procedural precision and optimize clinical outcomes. Despite substantial progress in developing intelligent perception systems and adaptive control architectures, fundamental challenges persist in high-fidelity data acquisition, protocol standardization, and the establishment of robust validation frameworks. The capabilities demonstrated by contemporary systems in controlled environments demonstrate both the promising potential and the inherent limitations of existing technologies, highlighting, in particular, the necessity for continued refinement in system scalability, operational reliability, and safety validation protocols.
Given the trajectory of robot-assisted endovascular procedures, we are likely to see imminent breakthroughs in autonomous capabilities and clinical applications. The emergence of sophisticated deep learning architectures, coupled with advances in multi-modal sensing technologies and real-time computational processing, suggests that innovations in autonomous navigation and decision support systems are imminent. Critical research priorities include the establishment of comprehensive benchmarking standards, the integration of interpretable AI frameworks for clinical decision support, and the development of robust control strategies capable of generalization across heterogeneous patient anatomies and pathologies. The achievement of these objectives necessitates sustained interdisciplinary collaboration between clinical practitioners and technical experts, along with rigorous adherence to regulatory requirements and ethical considerations. As these technologies mature, emphasis must be placed on developing stringent validation protocols, implementing standardized training methodologies, and establishing clear frameworks for clinical translation, with the ultimate goal of more precise, more efficient, and more accessible endovascular procedures.

### v2

@@
 Emerging trends and specific areas of research, such as federated learning for medical data sharing, explainable AI for clinical decision support, and advanced human-robot collaboration paradigms, are also explored, offering insights into the future direction of this rapidly evolving field.
-Index Terms:
-robotic-assisted endovascular procedures, medical image analysis, reinforcement learning
-I Introduction
+1 Introduction
 Endovascular procedures have been subject to a paradigm shift in relation to the management of vascular disease with the advent of minimally invasive therapeutic strategies that significantly shorten patient recovery times and enhance clinical outcomes [1].
@@
 The structure of this review is presented in Figure 1(d).
-Section II explores intelligent perception techniques in endovascular procedures, emphasizing recent progress in vascular segmentation and intervention device tracking.
+Section 2 explores intelligent perception techniques in endovascular procedures, emphasizing recent progress in vascular segmentation and intervention device tracking.
 These advancements are essential for improving the precision of such interventions.
-Section III explores the integration of machine learning in robotic-assisted endovascular procedures, providing an in-depth analysis of its role in advancing intelligence and refining control for these complex operations.
-Section IV assesses the major challenges facing the field, encompassing the limited availability of high-quality coronary imaging data, the lack of a standardized evaluation framework, constraints on the scalability of perception and control methodologies, ethical and legal implications, and the complexities involved in incorporating these technologies into clinical workflows.
+Section 3 explores the integration of machine learning in robotic-assisted endovascular procedures, providing an in-depth analysis of its role in advancing intelligence and refining control for these complex operations.
+Section 4 assesses the major challenges facing the field, encompassing the limited availability of high-quality coronary imaging data, the lack of a standardized evaluation framework, constraints on the scalability of perception and control methodologies, ethical and legal implications, and the complexities involved in incorporating these technologies into clinical workflows.
 Figure 2: Chord diagram depicting the interplay between perception and control modules in robotic-assisted endovascular procedures.
@@
 Arrows denote directional relationships, with line thickness reflecting connection strength, illustrating their integrated roles in enhancing procedural outcomes.
-II Intelligent Perception Techniques in Endovascular Procedures
+2 Intelligent Perception Techniques in Endovascular Procedures
 Endovascular procedures critically rely on precise and robust scene understanding to ensure effective navigation and task execution [17, 18].
@@
 By integrating high-resolution, real-time DSA with detailed 3D anatomical information from CTA, cardiologists can achieve superior navigation and enhanced visualization of complex vascular structures, ultimately improving procedural accuracy and clinical outcomes [26].
-II-A Vascular Segmentation Approaches
+2.1 Vascular Segmentation Approaches
 Vascular segmentation algorithms are carefully designed to delineate the vascular network with high precision, enabling accurate localization of target lesions and supporting optimal device deployment.
@@
 These advancements can be broadly categorized into three groups: (1) supervised learning, whereby large annotated datasets are employed to train models for precise vessel delineation; (2) weakly-supervised learning, which aims to diminish the dependency on labeled data; and (3) clinical applications, specifically tailored to address the challenges encountered in multi-center and real-world clinical studies.
-II-A1 Supervised Learning Methods
+2.1.1 Supervised Learning Methods
 Traditional supervised approaches have been shown to be effective in vascular segmentation, particularly through the utilization of the UNet architecture and its various adaptations, as illustrated in Figure 4.
@@
 The images are reproduced from [37], [38], [39], [22], and [40].
-II-A2 Weakly-supervised Learning
+2.1.2 Weakly-supervised Learning
 To address the challenge posed by limited labeled data, researchers have explored weakly-supervised strategies, particularly through the generation of pseudo-labels.
@@
 Their method utilizes label-free diffusion models and contrastive learning to extract meaningful features from unlabeled data.
-II-A3 Advanced Clinical Applications
+2.1.3 Advanced Clinical Applications
 Unlike research that focuses solely on algorithm optimization, studies aimed at clinical applications emphasize solutions that prioritize reliability, adaptability, and consistency.
@@
 proposed a geometry-constrained cascaded network, integrating geometric priors and multi-stage refinement to achieve Dice scores of 0.778 and 0.895 on the ASOCA and CCA-200 datasets respectively [47].
-II-A4 Available Public Datasets
+2.1.4 Available Public Datasets
 The availability of large-scale, publicly accessible datasets for coronary vascular segmentation remains limited, posing challenges for reproducibility and standardized benchmarking.
 Due to clinical regulations and privacy concerns, most current research relies on private datasets, further constraining the accessibility of standardized resources for broader evaluation and comparison.
-However, several valuable public datasets have been developed, as summarized in Table I.
+However, several valuable public datasets have been developed, as summarized in Table 1.
 The XCAD dataset [20], intended for weakly-supervised learning, comprises 1621 training frames and 126 test angiograms.
@@
 CCA-200 [47] presents 200 CCTA images annotated by professional radiologists, while CADICA [42] offers over 6,000 invasive coronary angiography images, supporting unsupervised learning approaches.
-TABLE I: Publicly Available Datasets for Coronary Vessel Segmentation
+Table 1: Publicly Available Datasets for Coronary Vessel Segmentation
 Dataset Description
@@
 Note: This table summarizes publicly available datasets for vascular segmentation research, including their size and content.
-II-B Intervention Device Perception
+2.2 Intervention Device Perception
 In endovascular procedures, catheters and guidewires can be effectively modeled as continuum robots [51].
@@
 This structured approach aims to facilitate a deeper understanding of the current state-of-the-art and to highlight future research directions in the field of intelligent perception for endovascular procedures.
-TABLE II: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception
+Table 2: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception
 No. Authors Category Method and Key Features Performance Metrics Dataset
@@
 7 Thuy et al.
-[61] Coarse-to-fine strategy for coronary vessel segmentation; Key Features: (1) UNet for the main coronary artery extraction (2) the second UNet for coronary artery-extracted segmentation Dice: 0.7640 Private
+[61] Coarse‐to‐fine strategy for coronary vessel segmentation; Key Features: (1) UNet for the main coronary artery extraction (2) the second UNet for coronary artery‐extracted segmentation Dice: 0.7640 Private
+Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union.
+Key metrics are defined as follows: F1 score is the harmonic mean of precision and recall, ranging from 0 to 1 (max = 1); Dice score (DSC) measures overlap between the predicted output ($A$) and ground truth ($B$) as $DSC=\frac{2|A\cap B|}{|A|+|B|}$, ranging from 0 to 1 (max = 1); AUC is the area under the ROC curve, ranging from 0 to 1 (max = 1); IoU is the intersection over union between the predicted output ($A$) and ground truth ($B$), $IoU=\frac{|A\cap B|}{|A\cup B|}$, ranging from 0 to 1 (max = 1).
+Additional metrics (AP, Acc, MPE) follow original definitions, with MPE in px or mm.
+Table 3: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
+No. Authors Category Method and Key Features Performance Metrics Dataset
 8 Tian et al. [62] Hybrid deep learning framework; Key Features: (1) Integrated digital processing (2) Multi-stage refinement (3) Morphological post-processing Dice: 0.939 Private
@@
 13 Zhang et al. [33] Recursive centerline-aware network; Key Features: (1) Direction-aware feature learning (2) Recursive refinement (3) Ensemble prediction F1: 0.831 Private
-14 Navab et al.
-[34] Weakly-supervised Learning for coronary vessel segmentation Contrastive diffusion-based learning; Key Features: (1) UNet with channel attention module (2) Centerline auxiliary supervision module Dice: 0.8528 Private
-15 Zhang et al. [41] Weakly-supervised segmentation with self-paced learning; Key Features: (1) Noisy label handling (2) Curriculum learning (3) Progressive refinement Dice: 0.821 Private
-Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union.
-Key metrics are defined as follows: F1 score is the harmonic mean of precision and recall, ranging from 0 to 1 (max = 1); Dice score (DSC) measures overlap between the predicted output ($A$) and ground truth ($B$) as $DSC=\frac{2|A\cap B|}{|A|+|B|}$, ranging from 0 to 1 (max = 1); AUC is the area under the ROC curve, ranging from 0 to 1 (max = 1); IoU is the intersection over union between the predicted output ($A$) and ground truth ($B$), $IoU=\frac{|A\cap B|}{|A\cup B|}$, ranging from 0 to 1 (max = 1).
-Additional metrics (AP, Acc, MPE) follow original definitions, with MPE in px or mm.
-TABLE III: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
+14 Navab et al. [34] Contrastive diffusion-based learning; Key Features: (1) UNet with channel attention module (2) Centerline auxiliary supervision module Dice: 0.8528 Private
+Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
+Table 4: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
 No. Authors Category Method and Key Features Performance Metrics Dataset
+15 Zhang et al.
+[41] Weakly-supervised Learning for coronary vessel segmentation Weakly-supervised segmentation with self-paced learning; Key Features: (1) Noisy label handling (2) Curriculum learning (3) Progressive refinement Dice: 0.821 Private
 16 Qi et al. [65] Examinee-examiner dual network; Key Features: (1) Centerline-guided learning (2) Weak supervision strategy (3) Dual network interaction Dice: 0.762 Private
@@
 [36] Localize all the endpoints of target vessel segments; Key Features: (1) Coarse-to-fine keypoint detection network (2) Multi-branch feature aggregation (3) Frangi filter with coordinate decoding for boundary enhancement, PCK: 0.763 Private
+Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
+Table 5: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
+No. Authors Category Method and Key Features Performance Metrics Dataset
 22 Lee et al.
@@
 28 Yao et al. [25] Integrated path planning system; Key Features: (1) Heuristic path optimization (2) Clinical workflow integration (3) Real-time guidance F1: 0.755 Private
+Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
+Table 6: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
+No. Authors Category Method and Key Features Performance Metrics Dataset
 29 Shen et al. [72] Vascular segmentation for clinical use; Key Features: (1) UNet with bidirectional ConvLSTM (2) Clinical workflow integration F1: 0.879 Private
@@
 ASOCA [48], CCA-200
-Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
-TABLE IV: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
-No. Authors Category Method and Key Features Performance Metrics Dataset
 33 Fu et al. [73] Robust temporal foreground extraction; Key Features: (1) Temporal consistency modeling (2) Noise-resistant features (3) Adaptive thresholding F1: 0.801 Private
@@
 35 Lalinia et al. [74] Edge-guided tracking framework; Key Features: (1) Multi-scale edge detection (2) Tracking optimization (3) Clinical integration Dice: 0.818 Private
+Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
+Table 7: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
+No. Authors Category Method and Key Features Performance Metrics Dataset
 36 Zhou et al. [75] Guidewire Detection and Tracking Pyramid attention recurrent network; Key Features: (1) Real-time processing (2) Multi-scale attention (3) Temporal modeling Dice: 0.942 Private
@@
 42 Zhang et al. [37] Real-time intraoperative tracking; Key Features: (1) Online tracking (2) Low latency (3) Clinical workflow integration Dice: 0.899 Private
+Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
+Table 8: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
+No. Authors Category Method and Key Features Performance Metrics Dataset
 43 Zhang et al. [80] Jigsaw transformer framework; Key Features: (1) Background reverse attention (2) Jigsaw training strategy (3) Multi-wire handling F1: 0.899 (single)
@@
 45 Wen et al. [82] Sim-to-real transfer framework; Key Features: (1) Domain adaptation (2) Synthetic data utilization (3) Cross-domain generalization IoU: 0.660 Cardiac [38]
-46 Du et al. [16] Guidewire endpoint localization via pixel-adjacent relation; Key Features: (1) UNet segmentation (2) Eight-neighborhood search mean pixel errors: 2.02 Private
+46 Du et al. [16] Guidewire endpoint localization via pixel‐adjacent relation; Key Features: (1) UNet segmentation (2) Eight‐neighborhood search mean pixel errors: 2.02 Private
 47 Gherardini et al.
@@
 Note: Acc = Accuracy; Sens = Sensitivity; Spec = Specificity; IoU = Intersection over Union
-TABLE V: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
+Table 9: Comprehensive Review of Data-Driven Methods for Endovascular Procedure Scene Perception (Continued)
 No. Authors Category Method and Key Features Performance Metrics Dataset
@@
 Single/dual refers to single or dual guidewire scenarios.
-II-B1 Guidewire Detection and Tracking
+2.2.1 Guidewire Detection and Tracking
 Guidewire detection and tracking are essential components in minimally invasive endovascular procedures, facilitating precise navigation through vascular pathways prior to catheter insertion.
@@
 Their method achieved an IoU of 0.660 on the public Cardiac dataset [38], demonstrating the potential of leveraging synthetic data to enhance guidewire detection models.
-II-B2 Catheter Detection and Tracking
+2.2.2 Catheter Detection and Tracking
 Catheter detection and tracking constitute critical components in minimally invasive cardiovascular procedures, directly enhancing procedural accuracy, facilitating real-time guidance, and minimizing the risk of complications [94, 95].
@@
 Their method achieved a mean error of 0.96 mm, showcasing its potential for highly accurate catheter localization in clinical settings.
-II-C Multi-modal Registration and Integration in Cardiovascular Imaging
+2.3 Multi-modal Registration and Integration in Cardiovascular Imaging
 Multi-modal registration and integration are crucial in cardiovascular imaging, as they provide enhanced accuracy in diagnosis and treatment [98].
@@
 Such learning-based methods surpass traditional approaches by offering higher accuracy and adaptability, which are critical in dynamic and complex clinical environments.
-TABLE VI: Summary of Data-Driven Control Methods for Endovascular Intervention Robots
+Table 10: Summary of Data-Driven Control Methods for Endovascular Intervention Robots
 No. Authors (Year) Task Category Method Performance Metrics
@@
 12 Mei et al. (2024) [109] Guidewire Navigation (Autonomous Delivery) Actor-critic architecture - Hierarchical learning - Knowledge transfer - Success rate - Learning speed - Decision accuracy
+Table 11: Summary of Data-Driven Control Methods for Endovascular Intervention Robots (continued)
+No. Authors (Year) Task Category Method Performance Metrics
 13 Wang et al.
@@
 (2024) [112] Intra-vascular Navigation DRL-based navigation: - End-to-end learning - Vision-based control - Safety constraints - Task success rate - Navigation efficiency - Safety metrics
-TABLE VII: Summary of Data-Driven Control Methods for Endovascular Intervention Robots (continued)
-No. Authors (Year) Task Category Method Performance Metrics
 17 Scarponi et al.
@@
 With a mean absolute error of 2.85 mm, this approach exemplifies how reinforcement learning can effectively navigate the complex space of registration parameters while maintaining clinical accuracy.
-III Learning-Based Control for Robotic-Assisted Endovascular Procedures
+3 Learning-Based Control for Robotic-Assisted Endovascular Procedures
 Endovascular devices constitute a unique class of continuum robots, characterized by their inherent flexibility and adaptability in navigating complex vascular networks [118].
@@
 By reducing the reliance on real-world data and enabling efficient algorithm development, simulation platforms, coupled with Sim2Real advancements, are poised to transform the landscape of robotic-assisted endovascular procedures.
-IV Challenges
+4 Challenges
 Data-driven approaches in endovascular robotic systems offer considerable promise, yet several critical challenges persist that need to be addressed to fully harness their potential.
 These challenges lie within the technical, ethical, and systemic domains - areas in which substantial progress is necessary to advance key capabilities and clinical adoption.
-IV-A Underdevelopment of Transformer-Based Models and Foundational Vision Models for Endovascular Procedures
+4.1 Underdevelopment of Transformer-Based Models and Foundational Vision Models for Endovascular Procedures
 Despite significant advancements in the use of transformer-based architectures and large pre-trained vision models in medical imaging and robotics [130, 131], their integration into endovascular procedures remains largely unexplored.
@@
 To address this, future research should prioritize the creation of large, multi-institutional, and multi-modal datasets that capture the broad diversity of patient anatomies, procedural contexts, and pathological conditions, all of which are crucial for training robust and generalizable models.
-IV-B Scarcity of High-Quality Coronary Imaging Data
+4.2 Scarcity of High-Quality Coronary Imaging Data
 Despite the extreme importance of having high-quality reference data in coronary imaging, current literature reveals surprisingly little research on this.
@@
 Addressing these data-related challenges is fundamental to advancing AI applications in coronary imaging and, ultimately, improving cardiovascular care through AI.
-IV-C Lack of Unified Workflow Recognition for Endovascular Procedures
+4.3 Lack of Unified Workflow Recognition for Endovascular Procedures
 The absence of a unified approach to workflow recognition in endovascular procedures poses a critical challenge, significantly impeding the advancement of robotic-assisted procedures and the development of higher-level autonomy in medical robotics [142, 143, 144, 145].
@@
 By augmenting a robot’s ability to anticipate and adapt to evolving instrument needs across procedural stages, this technology substantially improves both efficiency and safety, representing a pivotal step toward achieving intelligent endovascular robotic systems capable of high-level decision-making and adaptive behavior in dynamic environments.
-IV-D Absence of a Standardized, Large-Scale Benchmark
+4.4 Absence of a Standardized, Large-Scale Benchmark
 Recent advancements in large-scale benchmarking, exemplified by the Behavior-1K dataset [149] developed by Fei-Fei Li’s team, have significantly enhanced performance evaluation in endovascular interventions (EIs) by capturing 1,000 diverse, real-world activities within controlled simulation environments.
-The work deemed "Best Paper at ICRA 2024" also extended robotic learning through the introduction of standardized, high-quality datasets and models that support reproducibility and generalization across a wide range of robotic systems [150].
+The work deemed ”Best Paper at ICRA 2024” also extended robotic learning through the introduction of standardized, high-quality datasets and models that support reproducibility and generalization across a wide range of robotic systems [150].
 In contrast, endovascular robotics remains constrained by the scarcity of large-scale, domain-specific benchmarks that encapsulate the intricacies of human vasculature and the clinical complexity encountered in real-world settings.
@@
 By taking these steps, one could enhance both the technical sophistication, and clinical relevance, of endovascular robotic control, paving the way for better, safer, seamlessly integrated clinical procedures.
-IV-E Scalability of Perception and Learning-Based Control
+4.5 Scalability of Perception and Learning-Based Control
 Scalability remains a pivotal challenge for perception systems in endovascular procedures, given that these systems need to navigate the complex and dynamic vascular environment.
@@
 Additionally, transfer learning and domain adaptation approaches may aid in overcoming the challenge of generalizing control algorithms across different patient datasets.
-These strategies have the potential to substantially improve the flexibility and clinical applicability of endovascular robotic systems, ensuring their reliable performance across diverse patient populations, and ultimately leading to improved procedural outcomes and patient safety.
-IV-F Ethical and Legal Concerns
+These strategies have the potential to substantially improve the flexibility and clinical applicability of endovascular robotic systems, ensuring their reliable performance across diverse patient populations and ultimately leading to improved procedural outcomes and patient safety.
+4.6 Ethical and Legal Concerns
 The integration of autonomous robotic systems into medical procedures poses several ethical and legal challenges that must be addressed prior to their implementation.
@@
 Decision-making autonomy represents another critical issue.
-While AI-powered robotic systems can enhance precision, their "closed-box" nature, with opaque decision-making processes, raises concerns about clinician trust [157, 158].
+While AI-powered robotic systems can enhance precision, their ”closed-box” nature, with opaque decision-making processes, raises concerns about clinician trust [157, 158].
 Trustworthiness, defined as the reliability and predictability of a system, is crucial to ensure clinicians have confidence in the technology’s performance.
@@
 Efforts should be made to ensure equitable access to these technologies, ensuring they benefit a diverse range of patient populations.
-IV-G Integration with Existing Clinical Workflows
+4.7 Integration with Existing Clinical Workflows
 Beyond the technical challenges, the integration of robotic systems into existing clinical workflows poses a critical barrier to their widespread adoption.
 Although the potential of endovascular robotics to enhance precision, decrease procedural variability, and improve patient outcomes is evident, the transition from traditional manual methods to fully automated or semi-automated systems presents substantial challenges beyond purely technological aspects.
-A pivotal issue is the reluctance of clinical staff to embrace new systems, particularly when they are perceived as "black boxes" - autonomous systems with decision-making processes that are difficult to interpret or explain [158].
+A pivotal issue is the reluctance of clinical staff to embrace new systems, particularly when they are perceived as ”black boxes” - autonomous systems with decision-making processes that are difficult to interpret or explain [158].
 In endovascular procedures, in which precision and safety are paramount, the inability to fully comprehend how a robotic system arrives at its decisions can elicit significant levels of anxiety among medical professionals.
@@
 Without this educational investment, even the most advanced robotic technologies risk under-utilization in clinical practice.
-V Conclusion
+5 Conclusion
 The integration of data-driven methodologies into endovascular robotics signifies a paradigm shift in minimally invasive endovascular procedures.
@@
 As these technologies mature, emphasis must be placed on developing stringent validation protocols, implementing standardized training methodologies, and establishing clear frameworks for clinical translation, with the ultimate goal of more precise, more efficient, and more accessible endovascular procedures.
+Acknowledgement
+The authors would like to thank Mr.
+Mish Toszeghi from Queen Mary University of London for his meticulous proofreading of this manuscript.
+The authors also appreciate Mr.
+Haoyu Wang from Shanghai University of Engineering Science for his constructive discussions and valuable suggestions.
+This work is supported by the National Key Research and Development Program of China under Grant No.
+2023YFB4705200, the National Natural Science Foundation of China under Grant No.
+62273257, and the Open Project Fund of State Key Laboratory of Cardiovascular Diseases No.
+2024SKL-TJ002.
+Competing interests
+The authors declare no competing interests.